OpenStudy (firejay5):

Solve the inequalities and I will do the rest 1 - 4p/ 5 < or equal to 0.2 and 12(1/4 - n/3) < or equal to -6n

OpenStudy (firejay5):

they are fractions

OpenStudy (anonymous):

\[\frac{ 1-4p }{ 5 } \le0.2\]

OpenStudy (anonymous):

correct?

OpenStudy (firejay5):

yea I can't make it into a fraction

OpenStudy (anonymous):

first thing would be to multiply both sides of the inequality by 5 this will get rid of the fraction

OpenStudy (firejay5):

multiply 5 by each side

OpenStudy (firejay5):

okay I did that.

OpenStudy (anonymous):

ya so now you have \[1-4p \le1\]

OpenStudy (firejay5):

ok

OpenStudy (anonymous):

now subtract the 1 from both sides you now have \[-4p \le0\]

OpenStudy (anonymous):

now divide both sides by -4 now you have \[p \ge0\]

OpenStudy (firejay5):

that was the first one, there's 1 more I need answered

OpenStudy (anonymous):

k ill draw it out now

OpenStudy (firejay5):

k

OpenStudy (anonymous):

\[\frac{ 12 }{ 4 }-\frac{ 12n }{ 3 }<-6n\]

OpenStudy (anonymous):

i just distributed the 12 on the left side

OpenStudy (anonymous):

it reduces to \[3-4n <-6n\]

OpenStudy (anonymous):

thats supposed to be a less than or equal to sign sorry

OpenStudy (firejay5):

there's a line under inequality sign

OpenStudy (anonymous):

now add the 4n to both sides you now have \[3\le-2n\]

OpenStudy (anonymous):

now divide both sides by -2 your final answer: \[\frac{ -3 }{ 2 }\ge n\]

OpenStudy (firejay5):

what would it look opposite with N first

OpenStudy (anonymous):

\[n \le \frac{ -3 }{ 2 }\]

OpenStudy (firejay5):

is that correct

OpenStudy (anonymous):

i believe so lol

OpenStudy (firejay5):

how is this wrong: N > or equal to \[-\frac{ 3 }{ 2 }\]

OpenStudy (anonymous):

did you flip the inequality sign when you divided by -2

OpenStudy (firejay5):

yes, but I wanted n first

OpenStudy (anonymous):

well you can just write that after, when your solving the inequality you have to flip the inequality sign when you divide or multiply both sides by a negative number

OpenStudy (firejay5):

yes i know that

OpenStudy (anonymous):

what are you asking lol

OpenStudy (firejay5):

nevermind you're right I had to think it through

OpenStudy (anonymous):

alryte any questions let me know

OpenStudy (firejay5):

I need this once checked

OpenStudy (firejay5):

\[\frac{ d }{ 2 }\] > -4

OpenStudy (firejay5):

d > -8